Journalist Report – April 29th

Crew #: 335
Position: Crew Journalist
Report prepared by: Matteo Calore
Date: 29APR2026
Sol: 10

Report title: Driven by Mystery

Today we woke up as usual between 7 and 8. The day began in a dynamic way, with many discussions about the development of our projects. Afterwards, the film crew continued conducting interviews with the other crew members.
The interviews went very well, offering many interesting insights into personal lives and the motivations that brought us here, as well as what draws people toward space science and the mystery that surrounds the universe.
We often hear that one of the main motivations behind the study of the cosmos is the desire to understand how things work. The mystery of existence—why we are here, what makes everything what it is, and what keeps us moving and alive—remains largely unsolved. Yet it is precisely this mystery that drives many of us to search, to travel, to study, and to go to places we have never been before.

In the afternoon, we continued working on our projects. As we are approaching the end of our simulation, we are mostly wrapping things up, trying to understand what might still be missing and what can still be done.
The evening light was particularly beautiful, as was the view of the desert from our habitat. The gradual transitions of color across the clay hills—from red to white, to violet, to beige—and the scattered rocks that populate them make this experience truly unique.

GreenHab Report – April 29th

Report title: GreenHab Report

Crew #: 335

Position: GreenHab officer

Report prepared by: Charlotte Brugge

Date: 289R2026

Sol: day 9

Environmental control (fan & heater): Automated

Average temperatures (last 24h): 90.45

Maximum temperature (last 24h): 112.5

Minimum temperature (last 24h): 68.4

Hours of supplemental light: Automated

Daily water usage for crops: 0

Daily water usage for research and/or other purposes: 0

Water in the Tank (160 gal useful capacity): 86.73

Time(s) of watering for crops: N/A

Narrative: none, tomorrow demolition will begin

Harvest: (include which crop and mass in grams) 0

Support/supplies needed:0

Future GreenHab Improvement Observations: None noticed today

Sol Summary – April 28th

Report Title: Sol Summary Report
Crew #: 335
Position: Crew Commander
Report Prepared by: David Laude
Date: 28 Apr 2026
Sol: 9

Summary Title: Off to Green Mars View We Go

Mission Status: Mostly in good shape with continuing daily progress.

Sol Activity Summary: We were scheduled for EVA to Sea of Shells with four of us, but decided it best not to leave one ill crew member alone today and so instead, with approval of Mission Support, went with three to the stunning Green Mars View area where we did not need a fourth person for filming or Dave’s project. We stopped three times in the area along Cow Dung Rd. for filming near the View. Even with a late start and longer drive we got the job done and returned just before the scheduled end time. The rest of the day seemed so slow with the crew working on various projects. As usual, a nice lunch and dinner were prepared by our European crew mates.

Journalist Report – April 28th

Journalist Report

Crew #: 335

Position: Crew Journalist

Report Prepared by: Matteo Calore

Date: 28APR2026

Sol: 9

Report Title: Finding Our Rhythm

Today was a dynamic and demanding day. Some crew members went out for a long EVA to Green Mars View. The extended rover drive offered particularly striking and varied landscapes, ranging from clay-dense hills to stretches of desert sand punctuated by flowers of extraordinary beauty—vivid in color and twisted in form.

These flowers have an almost animal-like presence, and it feels remarkable that they can emerge with such strength from a terrain where water seems absent. It is astonishing how nature adapts to such different environments, continuously transforming and re-emerging in forms that are both diverse and somehow familiar.

The canyon at Green Mars View opens onto the valley where the Devil River flows. We found ourselves wondering how this river came to be given such a name. Despite its harsh surroundings, it carries life through the aridity, and in that sense, it seems to have little that is truly demonic.

While part of the crew was on EVA, the others continued working on their projects at the station. Charlotte transformed collected soil into clay and is now creating unique objects to be fired in the Science Dome’s microwave kiln. Kat continues her observations of the stars.

The day continued with interviews and various activities, and concluded with a generous meal, eagerly consumed by the crew of this small spaceship.

EVA Report – April 28th

EVA Report
Crew #: 335
Position: Elisa Strinna (Crew Artist)
EVA #: 12
Date: 28APR2026
Sol: 9

Purpose of EVA: To visit the Green Mars View. Crew members Elisa Strinna and Matteo Calore aim to film the area, Commander David Laude aim to realise his photographic project.

Weather Conditions: Overall good conditions, sunny with temperatures between 53°-64° F

Start Time: 10:20

End Time: 13:50

Narrative: After deciding to postpone our visit to the Sea of Shells to tomorrow because of the health conditions of one of the crew members we decided to head towards the Green Mars View to document the amazing view. We left a bit later then scheduled, overall the EVA went very well, not difficulties on the way, the landscape was very pleasant.

Destination: Green Mars View

Coordinates (use UTM WSG 84) : MDRS : 4251000N, 518000WGreen Mars View : 4257000N, 520500W

Route: Cow Dung Road

Mode of Travel: Rover + Walking

EVA Participants: David Laude (Commander) Elisa Strinna (Crew Artist), Matteo Calore (Crew HSO / Journalist)

Astronomy Report – April 28th

[category  astronomy-report]

Report title: Astronomy Report

Crew #: 335

Position: Engineer

Report prepared by: Katherine Berry

Date: 28Apr2026

Sol: 9

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: MDRS-WF

Objects to be Imaged this Evening: None, bad weather.

Images submitted with this report: None, images in process.

Problems Encountered: camera error for the MDRS–WF

MUSK OBSERVATORY

Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A

Operations Report – April 28th

Crew #: 335

Position: Engineer

Report prepared by: Katherine Berry

Date: 28Apr2026

Sol: 9

Non-nominal systems: None

Notes on non-nominal systems: None

ROVERS

Spirit rover used:

Hours: 309.8

Beginning charge: 100%

Ending charge: 34%

Currently charging: Yes

Opportunity rover:

Hours: 273.2

Beginning charge: 100%

Ending charge: 31%

Currently charging: Yes

Curiosity rover used:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Perseverance rover used:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

General notes on rovers: None

Summary of Hab operations:

Water Use (see notes): 20.53

Main tank level (remaining gallons, see notes): 198 gal

Main water tank pipe heater (ON or OFF): OFF

Main water tank heater (ON or OFF) OFF (was removed)

Toilet tank emptied (NO or YES): NO

Summary of internet: Connection was used for research, accessing, and working on homework, sending reports, video editing and listening to music.

Summary of suits and radios: None

Summary of GreenHab operations: (use information from the GreenHab report)

Water use: 2.88 gallons

Heater (ON or OFF): Auto

Supplemental light (hours of operation): Auto

Harvest (name, weight in grams): 0g

Summary of ScienceDome operations: Used to film creative projects, drying clay and manipulating clay

Dual split (Heat or AC, On or Off or Automatic): Automatic

Summary of RAM operations: one interview was held in the RAM.

Summary of any observatory issues: N/A

Summary of health and safety issues: Crew member that fell down the stairs is stable and has been up more.

Questions, concerns and requests to Mission Support: None

GreenHab Report – April 28th

Report title: GreenHab Report

Crew #: 335

Position: GreenHab officer

Report prepared by: Charlotte Brugge

Date: 28PR2026

Sol: day 8

Environmental control (fan & heater): Automated

Average temperatures (last 24h): 90.45

Maximum temperature (last 24h): 112.5

Minimum temperature (last 24h): 68.4

Hours of supplemental light: Automated

Daily water usage for crops: 2.88 gallons

Daily water usage for research and/or other purposes: 0

Water in the Tank (160 gal useful capacity): 86.73

Time(s) of watering for crops: 17.00

Narrative: watered and checked the plants.

Harvest: (include which crop and mass in grams) 0

Support/supplies needed:0

Future GreenHab Improvement Observations: None noticed today